OpenStudy (anonymous):

Solve: |2t + 5| = 3 A. {1, 4} B. {-1, -4} C. {-1, 4} D. {1, -4}

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

|2t + 5| = 3 turns into 2t+5 = 3 or 2t+5 = -3

jimthompson5910 (jim_thompson5910):

solve each equation for t

OpenStudy (anonymous):

2t when i dont understand who they got 3 or -3 :(

OpenStudy (anonymous):

well * how*

OpenStudy (anonymous):

wait i get it is the answer B?

jimthompson5910 (jim_thompson5910):

2t+5 = 3 or 2t+5 = -3 2t = 3-5 or 2t = -3-5 2t = -2 or 2t = -8 t = ?? or t = ??

OpenStudy (anonymous):

is the answer B ?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

:)